# Terawatt-level three-stage pulse compression for all-attosecond pump-probe spectroscopy

### v1

@@ Abstract
In conclusion, we have demonstrated a simple scheme for the compression of multi-mJ pulses down to 1.5 optical cycles. The high stability of the pulse duration and the pulse energy makes this approach ideal for the generation of intense attosecond pulses, which in our laboratory are now available as a turn-key system that is operational within a few minutes. The development of the pulse compression scheme has allowed us to perform APAPS in Ar and Ne that compares favorably to most previous APAPS demonstrations. This paves the way to performing even more demanding experiments such as all-attosecond transient absorption spectroscopy (AATAS). In the future, our approach may be scaled both to lower pulse energies (to reduce the required space), and to higher pulse energies and multi-TW powers. The required space in the latter case is compatible with the space that is available at large-scale facilities Kühn et al. (2017); Hort et al. (2019) and in university laboratories Rudawski et al. (2013); Wang et al. (2018) where suitable multi-TW lasers are available.

### v2

@@ 4 Conclusions and outlook
In conclusion, we have demonstrated a simple scheme for the compression of multi-mJ pulses down to 1.5 optical cycles. The high stability of the pulse duration and the pulse energy makes this approach ideal for the generation of intense attosecond pulses, which in our laboratory are now available as a turn-key system that is operational within a few minutes. The development of the pulse compression scheme has allowed us to perform APAPS in Ar and Ne with an unprecedented quality of the obtained data. This paves the way to performing even more demanding experiments such as all-attosecond transient absorption spectroscopy (AATAS). In the future, our approach may be scaled both to lower pulse energies (to reduce the required space), and to higher pulse energies and multi-TW powers. The required space in the latter case is compatible with the space that is available at large-scale facilities [34, 35] and in university laboratories [36, 37] where suitable multi-TW lasers are available.
@@ Acknowledgments
Disclosures
The authors declare no conflicts of interest.
Data availability statement
Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.
Supplemental document
See Supplement 1 for supporting content.